A wire of length 4.0 m with a cross-sectional

area of 3.2x10–6 m2 has a resistance of 0.065 Ω.

Determine the resistivity of the wire.



1
Expert's answer
2022-03-28T14:23:57-0400

The resistance of a wire

"R=\rho\frac{l}{A}"

Hence, the resistivity of a material

"\rho=\frac{RA}{l}=\frac{0.065*3.2*10^{-6}}{4.0}=5.2*10^{-8}\:\rm \Omega\cdot m"